# references.bib

### v1

@@ III Method
4:    if $t\text{ mod }N_{\text{inference}}=0$ then $\triangleright$ inference time step.
5:         $p_{t},I_{t},f_{t-T_{o}:t},q_{t}\leftarrow\textbf{{agent}}.\text{perception};$
6:         $\phi,a_{t:t+T_{a}}\leftarrow\text{FoAR}(p_{t},f_{t-T_{o}:t},I_{t});$
@@
20:    $\textbf{{agent}}.\text{execute}(a_{t});$triangleright execute the end-effector action.
@@ III-A Preliminary
Given an observation $p_{t}\in\mathbb{R}^{N_{t}\times 6}$ at current timestep $t$, RISE [rise] $\pi(p_{t})=a_{t:t+T_{a}}$ learns a direct mapping from the current observation to future robot actions over a horizon of $T_{a}$. Building upon RISE, our proposed force-aware policy, FoAR, incorporates high-frequency force/torque observations $f_{t-T_{o}:t}\in\mathbb{R}^{T_{o}\times 6}$ over a historical horizon of $T_{o}$ as additional inputs. Notice that $T_{o}$ represents the history horizon for high-frequency force/torque data, typically sampled at about 100Hz, while $T_{a}$ denotes the action horizon for future predictions, which operates at a lower frequency like 10Hz.
@@ III-B Force-Aware Policy Design
Point Cloud Encoder. Following RISE [rise], we employ sparse 3D encoder [minkowski] with a shallow ResNet architecture [resnet] to process the point cloud $p_{t}\in\mathbb{R}^{N_{t}\times 6}$ into sparse point tokens $P_{t}\in\mathbb{R}^{N_{p}\times 512}$. A Transformer [transformer] with sparse point encodings [rise] is then applied to these point tokens to generate a scene feature $h^{s}_{t}\in\mathbb{R}^{512}$.
@@ IV Experiments
  • Q3: Does reactive control during deployment enhance the policy’s ability to perform contact-rich actions?
@@ IV-A Setup
RISE (force-concat) 0.475 100 65 - - - 0.524 100 80
FoAR (3D-cls) 0.175 40 35 - - - 0.270 95 40
@@
TABLE I: Evaluation Results of the Wiping and Peeling Tasks. ASR denotes the action success rate. For the Wiping (General) task, we only select the best-performing ablation variant on the Wiping task, RISE (force-token), for evaluation.
Figure 4: Qualitative Results of the Peeling Task.
@@
Baselines. We evaluate our proposed approach against five baseline methods, including the vision-based policy RISE [rise] and three ablation variants: (1) RISE (force-token): incorporates encoded force/torque information as additional tokens within the RISE transformer, akin to [vtt, seehearfeel, maniwav, octo]; (2) RISE (force-concat): directly concatenates the force feature with the vision feature for action generation; (3) FoAR (3D-cls): uses scene features $h_{t}^{s}$ directly in the future contact predictor, instead of a separate image encoder.
@@ IV-B Surface Force Control Tasks: Wiping and Peeling
FoAR significantly outperforms baselines in surface force control tasks by integrating force/torque information to enhance manipulation accuracy and contact consistency in surface force control tasks (Q1). We report the evaluation results for the Wiping, Wiping (General), and Peeling tasks in Table I. Our proposed method, FoAR, achieves the highest scores of 0.875, 0.850, and 0.756 for the Wiping, Wiping (General), and Peeling tasks, respectively, significantly outperforming all baseline and variant methods. FoAR attains 100% success rates in both grasping the tool and performing the contact-rich operations (wiping and peeling) in all tasks, demonstrating its ability to maintain continuous and precise contact regardless of grasp position of the tool (eraser and peeler). In contrast, the pure vision-based policy RISE struggles with these contact-rich operations due to the lack of force/torque feedback, leading to inaccurate position control, which reflects the difficulty in maintaining consistent contact stemming from absence of force/torque perceptions. The qualitative results of the Peeling task in Fig. 4 further support these findings, showcasing that FoAR achieves more consistent and effective performance compared to the baselines, which often result in partial peelings or failures.
@@ IV-C Instantaneous Force Impact Task: Chopping
IV-D Ablations on Reactive Control
Reactive control in deployment enables the policy to perform more precise contact-rich actions (Q3). To illustrate the importance of reactive control during policy deployment, we use the Wiping (General) task as an example. The results in Table III demonstrate that reactive control is essential for achieving optimal policy performance. Notably, our proposed reactive control relies on the predicted future contact probability from the policy, highlighting the effective co-design of the FoAR policy and the reactive control mechanism. Consequently, in all experiments, reactive control is applied to FoAR-based methods to ensure high performance.
Method Score $\uparrow$ ASR(%) $\uparrow$
Grasp Wipe
FoAR wo. Reactive Control 0.650 100 80
@@ IV-D Ablations on Reactive Control
TABLE III: Ablation Results of the Wiping (General) Task on Reactive Control.
@@ IV-E Robustness to Dynamic Disturbances
FoAR maintains consistent task performance under unexpected and dynamic environmental disturbances, demonstrating superior robustness and adaptability (Q4). As shown in Tab. IV, FoAR achieves scores of 0.800, 0.850, and 0.800 for the Rewrite, Move, and Rewrite + Move robustness evaluations, respectively, while maintaining 100% action success rates. It can adapt to newly written random figures in erased areas, adjust its strategy to changes in whiteboard position, and handle both challenges simultaneously. As a strong vision-based baseline, RISE also exhibits decent generalization ability and handles these variations without performance degradation [rise], but its overall performance is constrained by the absence of force/torque feedback. Built upon RISE, FoAR successfully inherits the strong generalization ability, consistently detecting and responding to dynamic environmental changes while maintaining high performances. Its effective integration of force/torque information elevates performance to a higher level compared to RISE. Conversely, RISE (force-token) struggles to handle such complex scenarios and experiences a performance drop. We hypothesize that the unexpected disturbances force the policy to return to non-contact phases, requiring action sequence re-generations. The inherent noise in force/torque data during these phases further exacerbates errors in the re-generated sequences, hindering its effectiveness.
@@ V Conclusion
In this paper, we propose FoAR, a force-aware reactive policy tailored for contact-rich robotic manipulation. By introducing a future contact predictor, the policy enables effective contact-guided feature fusion between force/torque and visual information, dynamically balancing the contribution of each modality based on future contact probability. This design not only enhances precision during contact phases but also maintains strong performance in non-contact phases. Additionally, the future contact probability further guides the reactive control strategy, improving policy performance even with simple position control. Extensive experiments demonstrate the superior performance of FoAR in contact-rich tasks that require sustained and precise contact, such as wiping, peeling, and chopping. In the future, we plan to integrate advanced control strategies, such as compliance control and hybrid force/position control, into the FoAR policy to further enhance its performance. We also aim to extend this approach to dual-arm robots or humanoid robots for more complex contact-rich manipulation tasks.

### v2

@@ III Method
4:    if $t\text{ mod }N_{\text{inference}}=0$ then $\triangleright$ at the inference step.
5:         $p_{t},I_{t},f_{t-T_{o}:t},q_{t}\leftarrow\textbf{{agent}}.\text{perception};$$\triangleright$ perception.
6:         $\phi,a_{t:t+T_{a}}\leftarrow\text{FoAR}(p_{t},f_{t-T_{o}:t},I_{t});$triangleright inference.
@@
20:    $\textbf{{agent}}.\text{execute}(a_{t});$triangleright retrieve and execute the action.
@@ III-A Preliminary
Given an observation $p_{t}\in\mathbb{R}^{N_{t}\times 6}$ with $N_{t}$ points extracted from RGB-D image at current timestep $t$, RISE [rise] $\pi(p_{t})=a_{t:t+T_{a}}$ learns a direct mapping from the current observation to future robot actions over a horizon of $T_{a}$. Building upon RISE, our proposed force-aware policy, FoAR, incorporates high-frequency force/torque observations $f_{t-T_{o}:t}\in\mathbb{R}^{T_{o}\times 6}$ over a historical horizon of $T_{o}$ as additional inputs. Notice that $T_{o}$ represents the history horizon for high-frequency force/torque data, typically sampled at about 100Hz, while $T_{a}$ denotes the action horizon for future predictions, which operates at a lower frequency like 10Hz.
@@ III-B Force-Aware Policy Design
Point Cloud Encoder. Following RISE [rise], we employ sparse 3D encoder [minkowski] with a shallow ResNet architecture [resnet] to process the point cloud $p_{t}\in\mathbb{R}^{N_{t}\times 6}$ into sparse point tokens $P_{t}\in\mathbb{R}^{N_{p}\times 512}$, where $N_{p}$ represents the number of sparse point tokens after processing. A Transformer [transformer] with sparse point encodings [rise] is then applied to these point tokens to generate a scene feature $h^{s}_{t}\in\mathbb{R}^{512}$.
@@ IV Experiments
  • Q3: Can contact predictor, reactive control, and high-frequency force/torque sensing enhance the policy’s ability to perform contact-rich operations?
@@ IV-A Setup
RISE (force-concat) 0.475 100 65 0.675 100 95 0.524 100 80
FoAR (3D-cls) 0.175 40 35 0.200 40 40 0.270 95 40
@@
TABLE I: Evaluation Results of the Wiping and Peeling Tasks. ASR denotes the action success rate, measuring the success rates of the robot in executing certain actions, regardless of the quality of the actions.
Figure 4: Qualitative Results of the Peeling Task. The green background indicates the robot fully peels the cucumber, while yellow and red backgrounds represent partial peeling and peeling failure, respectively.
@@
Baselines. We evaluate our proposed approach against six baseline methods, including the vision-based policy ACT [act], Diffusion Policy [diffusionpolicy], and RISE [rise], as well as three ablation variants: (1) RISE (force-token): incorporates encoded force/torque information as additional tokens within the RISE transformer, akin to [vtt, seehearfeel, maniwav, octo]; (2) RISE (force-concat): directly concatenates the force feature with the vision feature for action generation; (3) FoAR (3D-cls): uses scene features $h_{t}^{s}$ directly in the future contact predictor, instead of a separate image encoder.
@@ IV-B Surface Force Control Tasks: Wiping and Peeling
FoAR significantly outperforms baselines in surface force control tasks by integrating force/torque information to enhance manipulation accuracy and contact consistency in surface force control tasks (Q1). We report the evaluation results for the Wiping, Wiping (General), and Peeling tasks in Table I. Our FoAR policy achieves the highest scores of 0.875, 0.850, and 0.756 for the Wiping, Wiping (General), and Peeling tasks, respectively, significantly outperforming all baselines and variants. FoAR attains 100% success rates in both grasping the tool and performing contact-rich operations (wiping and peeling) in all tasks, demonstrating its ability to maintain continuous and precise contact regardless of the grasp position of the tool (eraser and peeler). In contrast, the pure vision-based policy RISE struggles with these contact-rich operations due to the lack of force/torque feedback, leading to inaccurate position control, which reflects the difficulty in maintaining consistent contact stemming from the absence of force/torque perceptions. The qualitative results of the Peeling task in Fig. 4 further support these findings, showcasing that FoAR achieves more consistent and effective performance compared to the baselines, which often result in partial peelings or failures.
@@ IV-C Instantaneous Force Impact Task: Chopping
IV-D Ablations
Designing contact predictor, applying reactive control, and integrating high-frequency force/torque sensing enable the policy to perform more precise contact-rich manipulations (Q3). To illustrate the importance of these design, we take the Wiping task as an example. By comparing with a policy variant using contact detection instead of future contact prediction, we observe that the detection variant occasionally fails to make contact, indicating that contact detection may lead to delayed or even failed contact. Results in Tab. III also demonstrate that reactive control is essential for optimal policy performance. Notably, our proposed reactive control relies on the predicted future contact probability from the future contact predictor in the policy, highlighting the effective co-design of the FoAR policy and the reactive control mechanism. Additionally, relatively high-frequency (100Hz) force/torque sensing helps policy achieve better performance. Although there is a trade-off between high-frequency force/torque sensing and computational efficiency, we found 100 Hz as a balanced compromise.
Design Choice Score $\uparrow$ ASR (%) $\uparrow$
F/T Freq. w. Predictor w. Reactive Grasp Wipe
100Hz 0.650 100 85
100Hz 0.650 100 80
2Hz 0.625 100 85
10Hz 0.800 100 100
@@ IV-D Ablations
TABLE III: Ablation Results of the Wiping Task on Several Design Choices. We ablate our design choices on contact predictor, reactive control, and high-frequency force/torque sensing.
@@ IV-E Robustness to Dynamic Disturbances
FoAR maintains consistent task performance under unexpected and dynamic environmental disturbances, demonstrating superior robustness and adaptability (Q4). As shown in Tab. IV, FoAR successfully maintains consistent performance across all robustness evaluations, adapting to several dynamic environmental disturbances. While RISE also demonstrates strong generalization ability [rise], its performance is limited by the absence of force/torque feedback. Built upon RISE, FoAR inherits this generalization ability while leveraging force/torque integration to achieve superior performance. In contrast, RISE (force-token) struggles in these complex scenarios, likely due to disturbances forcing the policy into non-contact phases, requiring action re-generation. Noise in force/torque data during these transitions further amplifies errors, hindering its effectiveness.
@@ V Conclusion
In this paper, we propose FoAR, a force-aware reactive policy tailored for contact-rich robotic manipulation. By introducing a future contact predictor, the policy enables effective contact-guided feature fusion between force/torque and visual information, dynamically balancing the contribution of each modality based on future contact probability. This design not only enhances precision during contact phases but also maintains strong performance in non-contact phases. Additionally, the future contact probability further guides the reactive control strategy, improving policy performance even with simple position control. Extensive experiments demonstrate the superior performance of FoAR in contact-rich tasks that require sustained and precise contact, such as wiping, peeling, and chopping.
While FoAR demonstrates notable progress in contact-rich robotic manipulation, it has several limitations. The use of static force/torque thresholds, while effective in the evaluated tasks, may struggle in complex environments. Additionally, the current policy relies on the simple end-effector position control mechanism, limiting its adaptability. Future work will explore leveraging advanced control strategies, such as compliance control and hybrid force/position control, to enhance performance. Extending the FoAR policy to dual-arm robots, humanoid robots, and dexterous hands is another promising direction for enabling more complex and dexterous contact-rich manipulations.